An integration method that essentially involves using the chain rule in reverse is called…..
1
Expert's answer
2020-10-23T13:02:15-0400

"\displaystyle\textsf{Integration by Algebraic Substitution}\\

\textrm{Explanation:}\\

\textsf{For an Integral}\, S\\

S = \int f(x) \, \mathrm{d}x\\

\textsf{make the algebraic substitution}\\x = g(u)\\


g'(u) \mathrm{d}u = \mathrm{d}x\\


\therefore S = \int f(g(u)) g'(u)\mathrm{d}u \\


\textbf{\textsf{Example}}\\

\textsf{Integrate}\, \cos(r^2) r \mathrm{d}r\\

I = \int \cos(r^2) r \mathrm{d}r\\

\textsf{Let}\, u = r^2, \mathrm{d}u = 2r \mathrm{d}r \\

\begin{aligned}
I &= \int \frac{\cos(u)}{2r} r \mathrm{d}u \\&= \int \frac{\cos(u)}{2}du \\&= \frac{\sin(u)}{2} + C \\&= \frac{\sin(r^2)}{2} + C 
\end{aligned}"